OpenStudy (anonymous):

A ball has a radius of 5 cm. How many cubic cm are there?

Directrix (directrix):

What is the formula for the volume of a sphere? We need that to knock out this problem. @1j1

Directrix (directrix):

Recall that the formula to get the volume of a sphere is V = (4/3) × pi × r^3.

OpenStudy (anonymous):

I'm not sure how to calculate the formula.

geerky42 (geerky42):

Formula for the volume of a sphere: \(\large V = \dfrac{4}{3}\pi r^3\) Where \(V\) represents volume and \(r\) represents radius. All you have to do is plug in given value of radius then solve it and find the value of \(V\).

Directrix (directrix):